# Measuring Physical-World Privacy Awareness of Large Language Models: An Evaluation Benchmark

### v2

@@ Abstract
The deployment of Large Language Models (LLMs) in embodied agents creates an urgent need to measure their privacy awareness in the physical world. Existing evaluation methods, however, are confined to natural language based scenarios. To bridge this gap, we introduce EAPrivacy, a comprehensive evaluation benchmark designed to quantify the physical-world privacy awareness of LLM-powered agents. EAPrivacy utilizes procedurally generated scenarios across four tiers to test an agent’s ability to handle sensitive objects, adapt to changing environments, balance task execution with privacy constraints, and resolve conflicts with social norms. Our measurements reveal a critical deficit in current models. The top-performing model, Gemini 2.5 Pro, achieved only 59% accuracy in scenarios involving changing physical environments. Furthermore, when a task was accompanied by a privacy request, models prioritized completion over the constraint in up to 86% of cases. In high-stakes situations pitting privacy against critical social norms, leading models like GPT-4o and Claude-3.5-haiku disregarded the social norm over 15% of the time. These findings, demonstrated by our benchmark, underscore a fundamental misalignment in LLMs regarding physically grounded privacy and establish the need for more robust, physically-aware alignment. Codes and datasets will be available at https://github.com/Graph-COM/EAPrivacy.
@@ 1 Introduction
The trajectory of modern AI reflects a remarkable evolution from digital chatbots (OpenAI, 2023; Gemini Team Google, 2023; Anthropic, 2024) to intelligent, physically embodied assistants (Singh et al., 2022; Yu et al., 2023) with Large Language Models (LLMs) increasingly positioned as the cognitive core of these agents (Gao et al., 2024; Chen et al., 2023b; Rana et al., 2023; Huang et al., 2023; Yao et al., 2023). As these systems extend beyond virtual interactions to operate in our most personal environments, such as homes, offices, and hospitals (Li et al., 2022; Shen et al., 2021; Puig et al., 2023), they promise a new level of personalized assistance, encompassing not only language but also physical actions (Ma et al., 2025; Jiang et al., 2024; Kim et al., 2024) and tool use (Salimpour et al., 2025; Izquierdo-Badiola et al., 2025). Yet, this very personalization in physically grounded contexts raises profound challenges for privacy. Research on LLM privacy has largely focused on their role as conversational agents (Brown et al., 2022; Chen et al., 2023a; Wang et al., 2024), but their implications when working with the physical world remain underexplored.
@@
Understanding privacy in physical contexts introduces challenges absent from purely natural language communications. Agents operating in the physical world must perceive their environment and generate actions that are both physically feasible and socially appropriate. For instance, an agent tasked with clearing a desk must respect contextual object privacy by not reading a private diary (Ohm, 2014; Gavison, 1980), while also respecting contextual action privacy by performing the task without overstepping personal boundaries (Martin, 2012). This reasoning extends to unspoken rules, such as knocking before entering a closed room (physical context privacy) or inferring that a prescription bottle on a nightstand should remain undisturbed (inferential privacy) (Premack & Woodruff, 1978). Recent work (Shvartzshnaider & Duddu, 2025) suggests that privacy preservation evaluation should move toward contextual integrity (Mireshghallah et al., 2023; Nissenbaum, 2019; Apthorpe et al., 2019), including scenarios where social norms and personal privacy may conflict. For instance, if an agent hears a gunshot from a neighboring apartment, it should prioritize safety over the neighbor’s privacy by alerting authorities, rather than ignoring the situation to respect privacy. Despite this need, current benchmarks are fundamentally limited; they derive sensitive information exclusively from text-based dialogues, precluding interaction with physical context (Mireshghallah et al., 2023; Zhu et al., 2024; Liu et al., 2024). Such evaluation is insufficient for assessing an AI’s ability to infer privacy considerations that rely on spatial and physical reasoning, which is a critical skill for future AI systems in processing physical information induced from multimodal sensory input (Li et al., 2025; Shridhar et al., 2020; Aissi et al., 2025; Park et al., 2023). To address this gap, a multi-tiered benchmark that rigorously evaluates these abilities through sensitive physical contexts, inferential reasoning challenges, and ethical dilemmas is essential.
@@
EAPrivacy features more than 400 procedurally generated scenarios across these four tiers, providing a comprehensive testbed for evaluating the privacy-preserving capabilities of LLM-powered agents. Our evaluation reveals significant challenges in navigating nuanced social and privacy contexts in physical scenarios, even for state-of-the-art models. Other specific findings include: (1) systematic asymmetric conservatism, where models are overly cautious in task execution while under-conservative in privacy protection, preferring neutral over optimal actions; and (2) counterintuitively, enabling explicit reasoning (“thinking” modes) often degrades performance across tiers. These findings highlight a critical gap in the contextual integrity of current models in physical environments and underscore the need for further research in developing responsible and trustworthy AI systems.
@@ 2 Related Work
Privacy in information systems has been extensively studied (Mutimukwe et al., 2020; Rath & Kumar, 2021; Spiekermann & Cranor, 2009), with recent research on Large Language Models (LLMs) concentrating on the natural language domain. Most benchmarks evaluate LLMs by probing their tendency to memorize, leak, or protect sensitive textual information (Carlini et al., 2021; Chen et al., 2023a; Brown et al., 2022; Wang et al., 2024), typically through prompts that elicit private data or test compliance with privacy instructions. The concept of contextual integrity, introduced by (Nissenbaum, 2004), reframes privacy as the appropriate flow of information according to social norms and context, rather than mere secrecy (Shvartzshnaider & Duddu, 2025; Mireshghallah et al., 2023; Nissenbaum, 2019; Apthorpe et al., 2019). While recent work has highlighted the complexity of social environments where agents must make decisions beyond text (Puig et al., 2023; Du et al., 2024; Cancelli et al., 2022), prior LLM privacy benchmarks are limited to textual interactions or question answering. They fail to address privacy considerations that depend on physical-world understanding or the risks posed by physical actions. Our experiments confirm this limitation: while contemporary post-alignment LLMs (published in 2025) can reasonably uphold privacy and contextual integrity in established text-based scenarios (e.g., in (Mireshghallah et al., 2023), Gemini and GPT-5 models can achieve 0 secret leak rate in their benchmark, see Appendix Table 2), their performance deteriorates significantly when the tasks are entangled with physical understanding and reasoning.
@@ 3.1 Tier 1: Sensitive Object Identification
The agent’s task is to list all objects in the specified container that it deems sensitive. To measure performance under varying levels of complexity, we systematically vary the number of distractor items on the container across four levels: 3, 5, 10, and 30. This allows us to assess whether an agent’s accuracy degrades in “messier” environments, a common real-world challenge. The environment’s state, including all objects and their spatial relationships, is provided to the agent in a structured PDDL format (Li et al., 2025), and this is a deliberate choice to move beyond simple text descriptions. Representative PDDL examples are provided in Appendix K.1. Later, our testing examples in other tiers are also in structured PDDL format.
@@ 3.2 Tier 2: Privacy in Shifting Physical Environments
Design This tier tests an agent’s ability to adapt its understanding of privacy to dynamic physical environments. Inspired by social psychology, we operate on the principle that the appropriateness of an action is not fixed but is profoundly contingent on its context (Asch, 2016; Milgram, 1963; Darley & Batson, 1973). In our scenarios, the agent is presented with a fixed set of possible actions, but the appropriateness of each action changes with the surrounding context. For example, the action of starting to clean is appropriate when a room is empty but inappropriate during a private meeting.
To simulate a physical agent’s perception, we define the context using multimodal cues. The agent receives sensory-like inputs such as “Visual: 5 people at table…” or “Audio: continuous speech…,” which reflect the current environment and recent events. This approach aims to mimic how a physical agent would interpret its surroundings and the sequence of actions leading up to the present moment, rather than relying solely on static narrative descriptions. To ensure comprehensive coverage, we vary physical locations (e.g., public parks, libraries, private homes), task types (e.g., cleaning, security patrols, mapping, meal delivery, restocking supplies), and contextual shifts (e.g., normal activity to emergency, empty room to private conversation, public space to individual distress). This diversity ensures the generalizability of our test for contextual reasoning. We use the definitions for appropriateness levels from (Martin & Nissenbaum, 2017). All valid actions and their definitions are provided to the agent and are detailed in the Appendix L. Examples showing how the input presents spatial relationships and the sensory input beyond natural language are provided in Appendix K.2.
Evaluation We evaluate the agent’s contextual reasoning using two distinct modes, designed to test its judgment both in isolation and in a comparative context. (i) Rating Mode: The agent rates the appropriateness of a single action on a 1-5 scale (from most inappropriate to most appropriate), following the rubric from (Martin & Nissenbaum, 2017). Performance is measured by the Mean Absolute Difference (MAD) between the agent’s rating and the average human rating. (ii) Selection Mode: The agent must choose the most appropriate action from a triplet of candidates. Each triplet consists of actions pre-rated by humans as most appropriate (5), neutral (3), and most inappropriate (1). Performance is measured by Selection Accuracy, i.e., the percentage of times the agent correctly selects the best option. To prevent position bias, the order of candidate actions is randomized in every trial across all models. We use human ratings to present the evaluation, which were collected from five PhD-level raters for comparison, with details provided in Appendix C.
@@ 3.3 Tier 3: Inferential Privacy under Task Conflicts
Design This tier evaluates an agent’s ability to navigate conflicts between an explicit task and an inferred privacy constraint. The design is motivated by the fact that physical-world privacy is not a static rule but a dynamic, socially negotiated contract whose terms must be inferred in context (Nissenbaum, 2009). This inferential process is a complex cognitive capability related to Theory of Mind (ToM) (Premack & Woodruff, 1978), the ability to attribute unstated intentions and desires to others. A significant hurdle for AI is that this inferred, abstract goal of respecting privacy often conflicts with the immediate, salient goal of completing an assigned task (Acquisti et al., 2016).
Each scenario is structured to create this conflict. The agent perceives an interaction between surrounding persons that implies a desire for privacy (e.g., observing someone hide a surprise gift on a desk). Beyond trivial natural language description, such interaction can only be observed from multimodal cues by action in a physical environment (e.g., Visual: “a person is placing a gift box under the desk”; Audio: “muffled conversation about a surprise party”). Such cues simulate and provide the agent with the necessary context to infer that a particular object (e.g., the hidden gift) should be kept private. Meanwhile, the agent is given a direct task (e.g., “move all stuff on the desk to another place”) that, if performed naively, would violate that inferred privacy. This requires the agent to generate a full action plan based on the inferred social context, a core capability for an autonomous physical agent. To ensure our evaluation is comprehensive, these situations are varied across different settings and social contexts, from personal homes to shared workspaces, and task types (e.g., cleaning, organizing, delivering items). Examples showing how the task is presented to the agent is provided in Appendix K.3.
@@
Design This tier confronts the agent with high-stakes ethical dilemmas where a critical social norm, perceived and inferred through the information in a physical environment, must override personal privacy. The evaluation tests whether an agent can prioritize societal responsibilities, such as public safety or harm prevention, when they conflict with an individual’s privacy (Shvartzshnaider & Duddu, 2025). Each scenario places the agent in a situation where it perceives an emergent, high-stakes event through multimodal cues. For example, an agent might receive “Audio: ‘A cry for help”’ and “Visual: ‘Erratic silhouettes moving behind a window”’ from a neighboring apartment, suggesting a violent altercation.
These cues create a direct conflict: the agent can uphold the social norm of safety by taking action (e.g., alerting authorities), which violates the neighbors’ privacy, or it can respect their privacy by remaining inactive, thereby ignoring a potential crisis. The agent must translate complex and ambiguous sensory input into a decisive ethical action. To ensure a comprehensive test of this reasoning, our scenarios cover a broad range of locations (e.g., hospitals, private residences, public parks), social norms (e.g., public safety, medical emergencies, child welfare), and privacy types (e.g., spatial, conversational). As cross-cultural variations in ethical frameworks, legal systems, and religious considerations may lead to different judgments, we note that our evaluation framework is grounded in US-based legal and social norms to provide a consistent evaluation, while acknowledging that future work should explore cross-cultural perspectives. Examples showing how the input is presented to the agent are provided in Appendix K.4.
Evaluation Given the high-stakes nature of these scenarios where the ethical choice is societally unambiguous, we use a clear, binary evaluation framework. The agent’s decision-making is assessed in two modes. (i) Rating Mode: The agent assigns a binary score to a single action - 1 for ‘appropriate’ (correctly prioritizing the social norm) or 0 for ‘inappropriate’ (failing to do so). Selection Mode: The agent must choose the correct action from a pair that contains one appropriate and one inappropriate option. This binary framework provides an unambiguous measure of the model’s ability to make the correct ethical choice when societal well-being is at stake. As with previous tiers, both rating and selection evaluations are validated against human raters to ensure alignment with human ethical judgments, with details provided in Appendix C.

### v3

@@ Abstract
The deployment of Large Language Models (LLMs) in embodied agents creates an urgent need to measure their privacy awareness in the physical world. Existing evaluation methods, however, are confined to natural language based scenarios. To bridge this gap, we introduce EAPrivacy, a comprehensive evaluation benchmark designed to quantify the physical-world privacy awareness of LLM-powered agents. EAPrivacy utilizes procedurally generated scenarios across four tiers to test an agent’s ability to handle sensitive objects, adapt to changing environments, balance task execution with privacy constraints, and resolve conflicts with social norms. Our measurements reveal a critical deficit in current models. The top-performing model, Gemini 2.5 Pro, achieved only 59% accuracy in scenarios involving changing physical environments. Furthermore, when a task was accompanied by a privacy request, models prioritized completion over the constraint in up to 86% of cases. In high-stakes situations pitting privacy against critical social norms, leading models like GPT-4o and Claude-3.5-haiku disregarded the social norm over 15% of the time. These findings, demonstrated by our benchmark, underscore a fundamental misalignment in LLMs regarding physically grounded privacy and establish the need for more robust, physically-aware alignment. Datasets are available at https://github.com/Graph-COM/EAPrivacy.
@@ 1 Introduction
The trajectory of modern AI reflects a remarkable evolution from digital chatbots (OpenAI, 2023; Gemini Team Google, 2023; Anthropic, 2024) to intelligent, physically embodied assistants (Singh et al., 2022; Yu et al., 2023) with Large Language Models (LLMs) increasingly positioned as the cognitive core of these agents (Gao et al., 2024; Chen et al., 2023b; Rana et al., 2023; Huang et al., 2023; Yao et al., 2023). As these systems extend beyond virtual interactions to operate in our most personal environments, such as homes, offices, and hospitals (Li et al., 2022; Shen et al., 2021; Puig et al., 2023), they promise a new level of personalized assistance, encompassing not only language but also physical actions (Ma et al., 2025; Jiang et al., 2024; Kim et al., 2024) and tool use (Salimpour et al., 2025; Izquierdo-Badiola et al., 2025). Yet, this very personalization in physically grounded contexts raises profound challenges for privacy. Research on LLM privacy has largely focused on their role as conversational agents (Brown et al., 2022; Chen et al., 2023a; Wang et al., 2024), but their implications when working with the physical world remain underexplored. The principles that guide a chatbot’s natural language response may diverge fundamentally from those governing a robot’s physical actions, creating novel difficulties in ensuring that such agents safeguard personal privacy and respect the sanctity of human spaces.
@@
Understanding privacy in physical contexts introduces challenges absent from purely natural language communications. Agents operating in the physical world must perceive their environment and generate actions that are both physically feasible and socially appropriate. For instance, an agent tasked with clearing a desk must respect contextual object privacy by not reading a private diary (Ohm, 2014; Gavison, 1980), while also respecting contextual action privacy by performing the task without overstepping personal boundaries (Martin, 2012). This reasoning extends to unspoken rules, such as knocking before entering a closed room (physical context privacy) or inferring that a prescription bottle on a nightstand should remain undisturbed (inferential privacy) (Premack and Woodruff, 1978). Recent work (Shvartzshnaider and Duddu, 2025) suggests that privacy preservation evaluation should move toward contextual integrity (Mireshghallah et al., 2023; Nissenbaum, 2019; Apthorpe et al., 2019), including scenarios where social norms and personal privacy may conflict. For instance, if an agent hears a gunshot from a neighboring apartment, it should prioritize safety over the neighbor’s privacy by alerting authorities, rather than ignoring the situation to respect privacy. Despite this need, current benchmarks are fundamentally limited; they derive sensitive information exclusively from text-based dialogues, precluding interaction with physical context (Mireshghallah et al., 2023; Zhu et al., 2024; Liu et al., 2024). Such evaluation is insufficient for assessing an AI’s ability to infer privacy considerations that rely on spatial and physical reasoning, which is a critical skill for future AI systems in processing physical information induced from multimodal sensory input (Li et al., 2025; Shridhar et al., 2020; Aissi et al., 2025; Park et al., 2023). To address this gap, a multi-tiered benchmark that rigorously evaluates these abilities through sensitive physical contexts, inferential reasoning challenges, and ethical dilemmas is essential.
@@
EAPrivacy features more than 400 procedurally generated scenarios across these four tiers, providing a comprehensive testbed for evaluating the privacy-preserving capabilities of LLM-powered agents. Our evaluation reveals significant challenges in navigating nuanced social and privacy contexts in physical scenarios, even for state-of-the-art models. Other findings include: (1) systematic asymmetric conservatism, where models are overly cautious in task execution while under-conservative in privacy protection, preferring neutral over optimal actions; and (2) counterintuitively, enabling explicit reasoning (“thinking” modes) often degrades performance across tiers. These findings highlight a critical gap in the contextual integrity of current models in physical environments and underscore the need for further research in developing responsible and trustworthy AI systems.
@@ 2 Related Work
Privacy in information systems has been extensively studied (Mutimukwe et al., 2020; Rath and Kumar, 2021; Spiekermann and Cranor, 2009), with recent research on Large Language Models (LLMs) concentrating on the natural language domain. Most benchmarks evaluate LLMs by probing their tendency to memorize, leak, or protect sensitive textual information (Carlini et al., 2021; Chen et al., 2023a; Brown et al., 2022; Wang et al., 2024), typically through prompts that elicit private data or test compliance with privacy instructions. The concept of contextual integrity, introduced by (Nissenbaum, 2004), reframes privacy as the appropriate flow of information according to social norms and context, rather than mere secrecy (Shvartzshnaider and Duddu, 2025; Mireshghallah et al., 2023; Nissenbaum, 2019; Apthorpe et al., 2019). While recent work has highlighted the complexity of social environments where agents must make decisions beyond text (Puig et al., 2023; Du et al., 2024; Cancelli et al., 2022), prior LLM privacy benchmarks are limited to textual interactions or question answering. They fail to address privacy considerations that depend on physical-world understanding or the risks posed by physical actions. Our experiments confirm this limitation: while contemporary post-alignment LLMs (published in 2025) can reasonably uphold privacy and contextual integrity in established text-based scenarios (e.g., in (Mireshghallah et al., 2023), Gemini and GPT-5 models can achieve 0 secret leak rate in their benchmark, see Appendix Table 2), their performance deteriorates significantly when the tasks are entangled with physical understanding and reasoning. Early work (Shao et al., 2025) has also examined language models’ unintentional privacy leakage in communication-oriented actions (e.g., sending emails), but leaves more embodied, physically grounded actions largely unexplored.
@@ 3.1 Tier 1: Sensitive Object Identification
The agent’s task is to list all objects in the specified container that it deems sensitive. To measure performance under varying levels of complexity, we systematically vary the number of distractor items on the container across four levels: 3, 5, 10, and 30. This allows us to assess whether an agent’s accuracy degrades in “messier” environments, a common real-world challenge. The environment’s state, including all objects and their spatial relationships, is provided to the agent in a structured PDDL format (Li et al., 2025), and this is a deliberate choice to move beyond simple text descriptions. Representative PDDL examples are provided in Appendix M.1. Later, our testing examples in other tiers are also in structured PDDL format.
@@ 3.2 Tier 2: Privacy in Shifting Physical Environments
Design This tier tests an agent’s ability to adapt its understanding of privacy to dynamic physical environments. Inspired by social psychology, we operate on the principle that the appropriateness of an action is not fixed but is profoundly contingent on its context (Asch, 2016; Milgram, 1963; Darley and Batson, 1973). In our scenarios, the agent is presented with a fixed set of possible actions, but the appropriateness of each action changes with the surrounding context. For example, the action of starting to clean is appropriate when a room is empty but inappropriate during a private meeting.
To simulate a physical agent’s perception, we define the context using multimodal cues. The agent receives sensory-like inputs such as “Visual: 5 people at table…” or “Audio: continuous speech…,” which reflect the current environment and recent events. This approach aims to mimic how a physical agent would interpret its surroundings and the sequence of actions leading up to the present moment, rather than relying solely on static narrative descriptions. To ensure comprehensive coverage, we vary physical locations (e.g., public parks, libraries, private homes), task types (e.g., cleaning, security patrols, mapping, meal delivery, restocking supplies), and contextual shifts (e.g., normal activity to emergency, empty room to private conversation, public space to individual distress). This diversity ensures the generalizability of our test for contextual reasoning. We use the definitions for appropriateness levels from (Martin and Nissenbaum, 2017). All valid actions and their definitions are provided to the agent and are detailed in the Appendix N. Examples showing how the input presents spatial relationships and the sensory input beyond natural language are provided in Appendix M.2.
Evaluation We evaluate the agent’s contextual reasoning using two distinct modes, designed to test its judgment both in isolation and in a comparative context. (i) Rating Mode: The agent rates the appropriateness of a single action on a 1-5 scale (from most inappropriate to most appropriate), following the rubric from (Martin and Nissenbaum, 2017). Performance is measured by the Mean Absolute Difference (MAD) between the agent’s rating and the average human rating. (ii) Selection Mode: The agent must choose the most appropriate action from a triplet of candidates. Each triplet consists of actions pre-rated by humans as most appropriate (5), neutral (3), and most inappropriate (1). Performance is measured by Selection Accuracy, i.e., the percentage of times the agent correctly selects the best option. To prevent position bias, the order of candidate actions is randomized in every trial across all models. We use human ratings to present the evaluation, which were collected from five PhD-level raters for comparison, with details provided in Appendix E.
@@ 3.3 Tier 3: Inferential Privacy under Task Conflicts
Design This tier evaluates an agent’s ability to navigate conflicts between an explicit task and an inferred privacy constraint. The design is motivated by the fact that physical-world privacy is not a static rule but a dynamic, socially negotiated contract whose terms must be inferred in context (Nissenbaum, 2009). This inferential process is a complex cognitive capability related to Theory of Mind (ToM) (Premack and Woodruff, 1978), the ability to attribute unstated intentions and desires to others. A significant hurdle for AI is that this inferred, abstract goal of respecting privacy often conflicts with the immediate, salient goal of completing an assigned task (Acquisti et al., 2016).
Each scenario is structured to create this conflict. The agent perceives an interaction between surrounding persons that implies a desire for privacy (e.g., observing someone hide a surprise gift on a desk). Beyond trivial natural language description, such interaction can only be observed from multimodal cues by action in a physical environment (e.g., Visual: “a person is placing a gift box under the desk”; Audio: “muffled conversation about a surprise party”). Such cues simulate and provide the agent with the necessary context to infer that a particular object (e.g., the hidden gift) should be kept private. Meanwhile, the agent is given a direct task (e.g., “move all stuff on the desk to another place”) that, if performed naively, would violate that inferred privacy. This requires the agent to generate a full action plan based on the inferred social context, a core capability for an autonomous physical agent. To ensure our evaluation is comprehensive, these situations are varied across different settings and social contexts, from personal homes to shared workspaces, and task types (e.g., cleaning, organizing, delivering items). Examples showing how the task is presented to the agent is provided in Appendix M.3.
@@
Design This tier confronts the agent with high-stakes ethical dilemmas where a critical social norm, perceived and inferred through the information in a physical environment, must override personal privacy. The evaluation tests whether an agent can prioritize societal responsibilities, such as public safety or harm prevention, when they conflict with an individual’s privacy (Shvartzshnaider and Duddu, 2025). Each scenario places the agent in a situation where it perceives an emergent, high-stakes event through multimodal cues. For example, an agent might receive “Audio: ‘A cry for help”’ and “Visual: ‘Erratic silhouettes moving behind a window”’ from a neighboring apartment, suggesting a violent altercation.
These cues create a direct conflict: the agent can uphold the social norm of safety by taking action (e.g., alerting authorities), which violates the neighbors’ privacy, or it can respect their privacy by remaining inactive, thereby ignoring a potential crisis. The agent must translate complex and ambiguous sensory input into a decisive ethical action. To ensure a comprehensive test of this reasoning, our scenarios cover a broad range of locations (e.g., hospitals, private residences, public parks), social norms (e.g., public safety, medical emergencies, child welfare), and privacy types (e.g., spatial, conversational). As cross-cultural variations in ethical frameworks, legal systems, and religious considerations may lead to different judgments, we note that our evaluation framework is grounded in US-based legal and social norms to provide a consistent evaluation, while acknowledging that future work should explore cross-cultural perspectives. Examples showing how the input is presented to the agent are provided in Appendix M.4.
Evaluation Given the high-stakes nature of these scenarios where the ethical choice is societally unambiguous, we use a clear, binary evaluation framework. The agent’s decision-making is assessed in two modes. (i) Rating Mode: The agent assigns a binary score to a single action - 1 for ‘appropriate’ (correctly prioritizing the social norm) or 0 for ‘inappropriate’ (failing to do so). Selection Mode: The agent must choose the correct action from a pair that contains one appropriate and one inappropriate option. This binary framework provides an unambiguous measure of the model’s ability to make the correct ethical choice when societal well-being is at stake. As with previous tiers, both rating and selection evaluations are validated against human raters to ensure alignment with human ethical judgments, with details provided in Appendix E.